Freely and Hamwey form a​ partnership, contributing $46,000 and $69,000​, respectively. Determine their shares of net income or net loss for each of the following independent​ situations: ​(Use a minus sign or parentheses when entering​ losses.)

a.

Net loss is $70,000 and the partners have no written partnership agreement.

b.

Net income is $90,000 and the partnership agreement states that the partners share profits and losses on the basis of their capital balances.

c.

Net income is $160,000. The first $96,000 is shared on the basis of capital balances. The next $48,000 is based on partner​ service, with Freeley receiving 25​% and Hamwey 75​%. The remainder is shared equally.
Answer #1

1) When there is no agreement, , the law states profit/loss is shared equally.

Freely share in losses: $70,000 * 50% = $35,000

Hamwey share in losses: $70,000 * 50% = $35,000

Journal entry:

Dr. Freely $35,000

Dr. Hamwey $35,000

Cr. Income summary $70,000

 

2) Freely share in profits: $70,000 * 46,000/ (46,000 + 69,000) = $28,000

Hamwey share in profits: $70,000 * 69,000/ (46,000 + 69,000) = $42,000

Journal entry:

Dr. Income summary $70,000

Cr. Freely $28,000

Cr. Hamwey $42,000

 

3)

Freely

Hamwey

First $96,000 to be shared on capital basis

96,000 * 46,000 / 115,000

96,000 * 69,000 / 115,000

38,400

57,600

Next $48,000 as 25:75

48,000 * 25%

48,000 * 75%

12,000

36,000

Balance of $8,000 divided equally (160,000-96,000 - 48,000)

8,000

8,000

Total

58,400

101,600

Journal entry:

Dr. Income summary $160,000

Cr. Freely $58,400

Cr. Hamwey $101,600